# Generalized Linear Bandits: Almost Optimal Regret with One-Pass Update

### v1

@@ 1 Introduction
Stochastic multi-armed bandits (Robbins, 1952) represent a fundamental class of sequential decision-making problems where a learner interacts with environments by selecting actions (or arms) and receiving feedback in the form of rewards. The objective is to maximize the cumulative reward over time by balancing exploration and exploitation. In this paper, we focus on the contextual multi-armed bandit problem under the generalized linear model (GLM) framework (Filippi et al., 2010), where each action is associated with a contextual feature vector $\mathbf{x}\in\mathcal{X}_{t}\subset\mathbb{R}^{d}$, and the reward $r\in\mathbb{R}$ is generated according to a GLM satisfying $\mathbb{E}[r\,|\,\mathbf{x}]=\mu(\mathbf{x}^{\top}\theta_{*})$. Here, $\mu:\mathbb{R}\to\mathbb{R}$ is a non-linear link function in GLMs, and $\theta_{*}\in\mathbb{R}^{d}$ is an unknown parameter. Compared with the commonly studied linear case (Abbasi-Yadkori et al., 2011), the generalized linear bandit (GLB) framework allows for a richer class of reward distributions, including Gaussian, Bernoulli, Poisson, and Exponential distributions. This flexibility enables the modeling of various real-world applications, such as recommendation systems (Li et al., 2010) and personalized medicine (Tewari and Murphy, 2017), where the feedback is binary (Bernoulli) or count-based (Poisson) and inherently non-linear. Besides its practical appeal, the study of GLB canonically lays theoretical foundations for other sequential decision-making problems, such as function approximation in reinforcement learning (Wang et al., 2021), safe exploration (Wachi et al., 2021), and dynamic pricing (Chen et al., 2022).
[…]
@@
The non-linearity of the link function raises significant concerns regarding both computational and statistical efficiency in GLBs. A canonical solution to GLB is the GLM-UCB algorithm (Filippi et al., 2010), which estimates the underlying parameter $\theta_{*}$ via maximum likelihood estimation (MLE) and constructs an upper confidence bound (UCB) of the reward for arm selection. However, from a computational perspective, MLE is costly: it requires storing all historical data with $O(t)$ space complexity and solving an optimization problem with $O(t)$ time complexity at each round $t$. Besides, in terms of the statistical efficiency, the non-linearity of the link function introduces a notorious constant $\kappa=1/\inf_{\mathbf{x}\in\cup_{t=1}^{T}\mathcal{X}_{t},\theta\in\Theta}\mu^{\prime}(\mathbf{x}^{\top}\theta)$ into the regret bound of GLM-UCB (see Section 2 for details), where $\mu^{\prime}$ denotes the derivative of $\mu$. In several applications of GLBs, such as logistic bandits ($\mu(z)=1/(1+e^{-z})$) and Poisson bandits ($\mu(z)=e^{z}$), the $\kappa$ term can grow exponentially with the norm of the parameter $\lVert\theta_{*}\rVert_{2}$, severely affecting the theoretical performance.
@@
Over the past decade, extensive efforts have been devoted to enhancing the computational or statistical efficiency of GLBs. However, as summarized in Table 1, how to develop a jointly efficient method that achieves both low computation cost and strong statistical guarantees remains unclear. For GLBs, Jun et al. (2017) developed computationally efficient algorithms with one-pass update, but their regret bound scales linearly with the potentially large constant $\kappa$. More recently, by leveraging the self-concordance of the link function, Lee et al. (2024)Sawarni et al. (2024) and Liu et al. (2024) proposed statistically efficient methods that achieve improved dependence on $\kappa$; however, these approaches are still based on the MLE, which has high computation cost.
@@
Our Results. This paper proposes a jointly efficient algorithm for GLBs that achieves an improved regret bound in terms of $\kappa$ with constant time and space complexities per round as shown in Table 1. This advance roots in the construction of a tight confidence set for the online mirror descent (OMD) estimator used for performing the UCB-based exploration. We show that the OMD estimator, even though updated in a one-pass fashion, can still match the statistical efficiency of the MLE by carefully addressing the non-linearity of the link function. Here, one-pass refers to processing each data point only once, without storing past data. We also note that OMD-based online estimators have been used to develop jointly efficient algorithms in the logistic bandit setting (Faury et al., 2022; Zhang and Sugiyama, 2023; Lee and Oh, 2024), a special case of GLBs with Bernoulli rewards, their analyses of the confidence set rely heavily on the specific structure of the logistic link function $\mu(z)=1/(1+e^{-z})$, limiting their applicability to more general GLB models.
@@
$\mathcal{O}((\log T)^{\frac{3}{2}}\sqrt{T/\kappa_{*}})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$
@@
Technical Contribution.
Our main technical contribution is a new analysis of the OMD estimator to quantify its estimation error. The analysis is based on the concept of the mix loss, which has been used in full-information online learning to achieve fast-rate regret minimization (Vovk, 2001). Here, we show that it provides a natural way to bridge the loss gap between the OMD estimator and the true parameter $\theta_{*}$, thereby enabling the construction of tight confidence sets for bandit online learning. Our new analysis not only generalizes the OMD-based approach to the broader GLB setting but also improves upon the state-of-the-art for logistic bandits. As shown in Table 2, the jointly efficient method (Faury et al., 2022) requires $\mathcal{O}(\log t)$ time per round and an adaptive warm-up strategy to achieve optimal regret. More recent method (Zhang and Sugiyama, 2023) reduces the time complexity to $\mathcal{O}(1)$ but incurs an extra $\mathcal{O}(\sqrt{\log T})$ factor in the regret. In contrast, our refined analysis yields a tighter error bound for the OMD estimator, allowing our method to achieve improved regret and low computation cost without warm-up rounds. More details are provided in Section 4.
@@ 2 Preliminary
The GLB problem considers a $T$-round sequential interaction between a learner and the environment. At each round $t\in[T]\triangleq\{1,\dots,T\}$, the learner selects an action $X_{t}\in\mathcal{X}_{t}\subset\mathbb{R}^{d}$ from the feasible domain and then receives a stochastic reward $r_{t}\in\mathbb{R}$. We denote the learner’s action by $X_{t}$ to emphasize its stochastic nature, which may depend on past data captured by the filtration $\mathcal{F}_{t}=\sigma(X_{1},r_{1},\dots,X_{t-1},r_{t-1})$. In GLBs, conditioned on $\mathcal{F}_{t}$, the reward $r_{t}$ follows a canonical exponential family distribution with the natural parameter given by the linear model $z_{t}=X_{t}^{\top}\theta_{*}$.
@@
In the above, $\theta_{*}\in\mathbb{R}^{d}$ is a $d$-dimensional vector unknown to the learner. The function $h(r,\tau)$ is the base measure, which provides the intrinsic weighting of the variable $r$, while $m(z)$ is twice continuously differentiable function used for normalizing the distribution. $g:\mathbb{R}\rightarrow\mathbb{R}$ is the dispersion function controlling the variability of the distribution. It is known that the expectation and variance of the exponential family distribution can be calculated as $\mathbb{E}[r\>|\>z]=\mu(z)\triangleq m^{\prime}(z)$ and $\mbox{Var}(r\>|\>z)=g(\tau)m^{\prime\prime}(z)$ (Wainwright and Jordan, 2008, Proposition 3.1). Common examples of (1) include Gaussian, Bernoulli, Poisson, and Exponential distributions. The goal of the learner is to maximize the cumulative expected reward, which is equivalent to minimizing the regret,
@@
where $\mathbf{x}_{t,*}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}_{t}}\mu(\mathbf{x}^{\top}\theta_{*})$ is the optimal action. Besides, we have the following assumptions, which are standard boundness assumptions in the GLB literature (Filippi et al., 2010).
@@
The link function $\mu$ is twice differentiable over its feasible domain. Moreover, there exist constants $c_{\mu}>0$ and $C_{\mu}>0$ such that $c_{\mu}\leq\mu^{\prime}(z)\leq C_{\mu}$ for all $z\in[-S,S]$. Consequently, the function $m$ is strictly convex and $\mu$ is strictly increasing.
GLM-UCB Method and Potentially Large Constant.
The canonical algorithm for the GLB problem is GLM-UCB (Filippi et al., 2010), which resolves the exploration-exploitation trade-off with an upper‑confidence‑bound strategy (Agrawal, 1995). Under Assumptions 12 and an additional condition that the reward $r_{t}$ is non-negative and almost surely bounded for all $t\in[T]$, GLM-UCB achieves the regret of $\mathcal{O}\big{(}\kappa(\log T)^{\frac{3}{2}}\sqrt{T}\big{)}$, where the $\mathcal{O}(\cdot)$-notation is used to highlight the dependence on $\kappa$ and time horizon $T$. This bound is nearly optimal in terms of the horizon $T$ up to logarithmic factors. The bottleneck lies in its linear dependence on the constant $\kappa$, which is defined by
@@ 2.2 New Progress with Self-Concordance
The undesirable linear dependence on $\kappa$ has motivated the development of algorithms with improved theoretical guarantees. Recent work (Russac et al., 2021; Lee et al., 2024; Sawarni et al., 2024) showed that leveraging the self-concordance of the link function yields regret bounds with significantly reduced dependence on $\kappa$. Following this line, we also adopt the self-concordance assumption here.
@@
Assumption 3 holds for many widely used GLMs. For GLMs where the reward is almost surely bounded in $[0,R]$, the link function satisfies Assumption 3 with coefficient $R$ (Sawarni et al., 2024). For example, the Bernoulli distribution is $1$-self-concordant. Many unbounded GLMs also satisfy self-concordance, including the Gaussian distribution ($R=0$), Poisson distribution ($R=1$), and Exponential distribution ($R=0$). Leveraging self-concordance, Lee et al. (2024) and Sawarni et al. (2024) established improved regret bounds of order $\mathcal{O}(\log T\sqrt{T/\kappa_{*}})$. In these results, the potentially large constant $\kappa_{*}$ appears at the denominator, indicating that the non-linearity is beneficial to regret minimization. However, their methods incur still $\mathcal{O}(t)$ time and space complexities per round. Our goal is to design a method with low computation cost while maintaining strong regret guarantees.
@@
Our GLM assumptions are more general than the canonical GLM formulation introduced in Filippi et al. (2010) and later adopted in Sawarni et al. (2024). Besides Assumptions 1 and 2Filippi et al. (2010) further requires the rewards to be almost surely bounded, which automatically implies self-concordance and thus satisfies Assumption 3 as shown by Sawarni et al. (2024, Lemma 2.2). Beyond bounded distributions, our GLM formulation accommodates unbounded ones, such as Gaussian or Poisson. ◆
@@ 3.1 OFU Principle and Computational Challenge
Computational Challenge.
To the best of our knowledge, most existing OFU-based methods for GLBs rely on maximum likelihood estimation (MLE) to estimate $\theta_{*}$ and construct the confidence set. Specifically, the estimator is computed as
@@ 3.2 Jointly Efficient Method
~$\widetilde{\ell}_{t}(\theta)\triangleq\langle\nabla\ell_{t}(\theta_{t}),\theta-\theta_{t}\rangle+\frac{1}{2}\lVert\theta-\theta_{t}\rVert^{2}_{\nabla^{2}\ell_{t}(\theta_{t})}~{}~{}~{}\mbox{and}~{}~{}~{}H_{t}\triangleq\lambda I_{d}+\sum_{s=1}^{t-1}\nabla^{2}\ell_{s}(\theta_{s+1}).$ (4)
The above two components play important roles in achieving low computation cost while maintaining statistical efficiency. The loss function ~$\widetilde{\ell}_{t}(\theta)$ serves as a second-order approximation of the original loss, which preserves the curvature information of the current loss function while ensures that the resulting optimization problem can be solved efficiently. The local matrix can also be expressed as $H_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}\mu^{\prime}(X_{s}^{\top}\theta_{s+1})/g(\tau)\cdot X_{s}X_{s}^{\top}$, where $X_{s}\in\mathbb{R}^{d}$ is the action selected by the learner. The matrix explicitly captures the non-linearity of the link function by $\mu^{\prime}(X_{s}^{\top}\theta_{s+1})$ and retains the curvature information of historical loss functions until time $t-1$. Since the optimization problem (3) is quadratic optimization over an Euclidean ball, it can be solved with a computation cost of $\mathcal{O}(d^{3})$, independent of time $t$. Further details on solving (3) are provided in Appendix D.
@@
Let $\delta\in(0,1]$. Set the step size to $\eta=1+RS$ and the regularization parameter to $\lambda=2\max\{7d\eta R^{2},\max\{3\eta RS,1\}C_{\mu}/g(\tau)\}$. For each $t\in[T]$, define the confidence set as
@@
$\beta_{t}(\delta)=\sqrt{4\lambda S^{2}+2\eta\ln\Bigl{(}\frac{1}{\delta}\Bigr{)}+6d\eta^{2}\ln\Bigl{(}2+\frac{2C_{\mu}t}{\lambda g(\tau)}\Bigr{)}}=\mathcal{O}\left(SR\sqrt{d\Bigl{(}S^{2}R+\ln\Bigl{(}\frac{t}{\delta}\Bigr{)}\Bigr{)}}\right).$
Then, under Assumptions 1, 2, and 3, we have $\Pr\Bigl{[}\forall t\geq 1,\;\theta_{*}\in\mathcal{C}_{t}(\delta)\Bigr{]}\geq 1-\delta.$
Theorem 1 shows that an ellipsoidal confidence set $\mathcal{C}_{t}(\delta)$ can be constructed to quantify the uncertainty of the online estimator $\theta_{t}$ with both statistical and computational efficiency. From a computational perspective, constructing the confidence set relies only on the online estimator, which can be updated with $\mathcal{O}(1)$ time and space complexities. From a statistical view, the radius of the confidence set is independent of $\kappa$, which is crucial for achieving the improved regret bound, as detailed next.
@@
We note that the OMD has been used in the logistic bandits for constructing confidence sets (Faury et al., 2022; Zhang and Sugiyama, 2023; Lee and Oh, 2024). However, existing methods are not fully jointly efficient compared with our result. Specifically, Faury et al. (2022) required optimization over the original loss at each round, incurring an additional $\mathcal{O}(\log t)$ computation cost and relying on an adaptive warm-up strategy to maintain statistical efficiency. In later work, Zhang and Sugiyama (2023) and Lee and Oh (2024) achieved a constant per-round cost but their regret bounds have an extra $\mathcal{O}(\sqrt{\log t})$ multiplicative factor. More importantly, as we will discuss in detail in Section 4, the analyses in these works depend on the specific structure of the logistic model and do not naturally extend to the GLB setting. Our main contribution is to introduce the notion of mix loss into the confidence set analysis, which not only enables the application of OMD to the broader GLB framework but also improves both statistical and computational efficiency over previous methods for logistic bandits. ◆
Arm Selection and Regret Bound.
Based on the ellipsoidal confidence set, one can employ a variety of exploration strategies, not limited to OFU-based methods but also including randomized approaches such as Thompson sampling (Abeille and Lazaric, 2017) for action selection. Here, we adopt the classical OFU-based strategy, where the action $X_{t}$ is selected by solving the bilevel optimization problem $X_{t}=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{X}_{t}}\max_{\theta\in\mathcal{C}_{t}(\delta)}\mu(\mathbf{x}^{\top}\theta).$ Since $\mu$ is an increasing function and $\mathcal{C}_{t}(\delta)$ is an ellipsoid, the OFU-based action selection rule is equivalent to
@@
0:  Self-concordant constant $R$, Lipchitz constant $C_{\mu}$, parameter radius $S$, confidence level $\delta$.
@@
Let $\delta\in(0,1]$. Under Assumptions 1, 2, and 3, with probability at least $1-\delta$, Algorithm 1 with parameter $\eta=1+RS$ and $\lambda=2\max\{7d\eta R^{2},\frac{\max\{3\eta RS,1\}C_{\mu}}{g(\tau)}\}$ ensures

### v2

@@ 1 Introduction
Stochastic multi-armed bandits (Robbins, 1952) represent a fundamental class of sequential decision-making problems where a learner interacts with environments by selecting actions (or arms) and receiving feedback in the form of rewards. In this paper, we study the contextual multi-armed bandit problem under the framework of generalized linear models (GLMs). In this setting, each action is characterized by a contextual feature vector $\mathbf{x}\in\mathcal{X}_{t}\subset\mathbb{R}^{d}$, where the arm set $\mathcal{X}_{t}$ may vary over time. More specifically, the learning process can be seen as a $T$ round game between the learner and environments: at each round $t$, the learner selects an action $X_{t}\in\mathcal{X}_{t}$ and then observes a stochastic reward $r_{t}\in\mathbb{R}$ generated according to a GLM (see Definition 2.1). The goal of the learner is to maximize the cumulative expected reward obtained over the time horizon $T$. Under the GLM model, the expectation of the reward satisfies $\mathbb{E}[r_{t}\,|\,X_{t}]=\mu(X_{t}^{\top}\theta_{*})$, where $\mu:\mathbb{R}\to\mathbb{R}$ is a non-linear link determined by the GLM model and is known to the learner. The unknown part is the underlying parameter $\theta_{*}\in\mathbb{R}^{d}$, which needs to be estimated from the observed action-reward pairs.
Compared with the classical linear case (Abbasi-Yadkori et al., 2011), the generalized linear bandit (GLB) framework allows for a richer class of reward distributions, including Gaussian, Bernoulli, and Poisson distributions. This flexibility enables the modeling of various real-world tasks, such as recommendation systems (Li et al., 2010) and personalized medicine (Tewari and Murphy, 2017), where the feedback is binary (Bernoulli) or count-based (Poisson) and inherently non-linear. Besides its practical appeal, the study of GLB lays theoretical foundations for other sequential decision-making problems, such as function approximation in RL (Wang et al., 2021; Li et al., 2024), safe exploration (Wachi et al., 2021), and dynamic pricing (Chen et al., 2022; Xu and Wang, 2021).
Table 1: Comparison of regret guarantees and computational complexity per round for GLBs. Here, $\kappa_{*}=1/\big(\frac{1}{T}\sum_{t=1}^{T}\mu^{\prime}(\mathbf{x}_{t,*}^{\top}\theta_{*})\big)$ is the slope at the optimal action $\mathbf{x}_{t,*}=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{X}_{t}}\mu(\mathbf{x}^{\top}\theta_{*})$, with $\kappa_{*}\leq\kappa$ (see Section 2 for details). $\dagger$} indicates the amortized time complexity, i.e., average per-round cost over $T$ rounds.
@@
The non-linearity of the link function raises significant concerns regarding both computational and statistical efficiency in GLBs. A canonical solution to GLB is the GLM-UCB algorithm (Filippi et al., 2010), which belongs to the family of UCB-type methods (Agrawal, 1995; Auer, 2002). At each iteration $t\in[T]$, the algorithm estimates the true parameter $\theta_{*}$ using maximum likelihood estimation (MLE) based on the historical data $\{(\mathbf{x}_{s},r_{s})\}_{s=1}^{t-1}$ and yields an estimator $\theta_{t}$, which is further used for constructing the upper confidence bound for arm selection. As shown in Table 1, GLM-UCB achieves nearly optimal regret bound in terms of the dependence on $T$. However, its reliance on MLE incurs a computational burden: it requires storing all historical data with $\mathcal{O}(t)$ space complexity and solving an optimization problem with $\mathcal{O}(t)$ time complexity at each round $t$. Besides, in terms of the statistical efficiency, the non-linearity of the link function introduces a notorious constant $\kappa=1/\inf_{\mathbf{x}\in\cup_{t=1}^{T}\mathcal{X}_{t},\theta\in\Theta}\mu^{\prime}(\mathbf{x}^{\top}\theta)$ into the regret bound of GLM-UCB (see Section 2 for details), where $\mu^{\prime}$ denotes the derivative of $\mu$. In several applications of GLBs, such as logistic bandits ($\mu(z)=1/(1+e^{-z})$) and Poisson bandits ($\mu(z)=e^{z}$), the $\kappa$ term can grow exponentially with the norm of the parameter $\lVert\theta_{*}\rVert_{2}$, severely affecting the theoretical performance.
@@
Over the past decade, extensive efforts have been devoted to enhancing the computational or statistical efficiency of GLBs. However, as summarized in Table 1, how to develop a jointly efficient method that achieves both low computation cost and strong statistical guarantees remains unclear. For GLBs, Jun et al. (2017) developed computationally efficient algorithms with one-pass update, but their regret bound scales linearly with the potentially large constant $\kappa$. More recently, by leveraging the self-concordance of the loss, several works (Lee et al., 2024; Sawarni et al., 2024; Liu et al., 2024; Clerico et al., 2025) proposed statistically efficient methods that achieve improved dependence on $\kappa$; however, these approaches are still based on the MLE, which has high computation cost.
@@
Our Results. This paper proposes a jointly efficient algorithm for GLBs that achieves an improved regret bound in terms of $\kappa$ with constant time and space complexities per round as shown in Table 1. This advance roots in the construction of a tight confidence set for the online mirror descent (OMD) estimator used for performing the UCB-based exploration. We show that the OMD estimator, even though updated in a one-pass fashion, can still match the statistical efficiency of the MLE by carefully addressing the non-linearity of the link function. Here, “one-pass” refers to processing each data point only once, without storing past data. We also note that OMD-based online estimators have been used to develop jointly efficient algorithms in the logistic bandit setting (Faury et al., 2022; Zhang and Sugiyama, 2023; Lee and Oh, 2024), a special case of GLBs with Bernoulli rewards. However, their analyses of the confidence set rely heavily on the specific structure of the logistic link function $\mu(z)=1/(1+e^{-z})$, which limits their applicability to the more general GLB setting.
@@
$\mathcal{O}((\log T)^{3/2}\sqrt{T/\kappa_{*}})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$
@@
Technical Contribution. Our main technical contribution is a new analysis of the estimation error of the OMD estimator. The analysis is based on the concept of the mix loss, which has been used in full-information online learning to achieve fast-rate regret minimization (Vovk, 2001). Here, we show that it provides a natural way to bridge the gap between the OMD estimator and the true parameter $\theta_{*}$, thereby enabling the construction of tight confidence sets for bandit online learning. Our new analysis not only generalizes the OMD-based approach to the broader GLBs but also improves upon the state-of-the-art for logistic bandits. As shown in Table 2, the jointly efficient method (Faury et al., 2022) requires $\mathcal{O}(\log t)$ time per round and an adaptive warm-up strategy to achieve optimal regret. Zhang and Sugiyama (2023) reduces the time complexity to $\mathcal{O}(1)$ but incurs an extra $\mathcal{O}(\sqrt{\log T})$ factor in the regret. In contrast, our refined analysis yields a tighter error bound for the OMD estimator, allowing our method to achieve improved regret and low computation cost without warm-up rounds. Details are provided in Section 4.
We were made aware that mix-loss–based analyses have been independently developed in two very recent concurrent works for constructing tight confidence sets. Specifically, Kirschner et al. (2025) developed confidence sets based on the sequential likelihood ratios mixing technique, and Clerico et al. (2025) proposed several confidence sets for GLMs. While conceptually related, these works focus on the batch setting, where all historical data are repeatedly accessed, leading to substantial computational overhead. In contrast, our confidence set is based on the OMD estimator with a one-pass update. This difference leads to a distinct formulation of the mix loss and a tailored analysis to quantify its gap relative to the time-varying OMD estimator. Details are provided in Section 4.2.
@@ 2 Preliminary
The GLB problem considers a $T$-round sequential interaction between a learner and the environment. At each round $t\in[T]\triangleq\{1,\dots,T\}$, the learner selects an action $X_{t}\in\mathcal{X}_{t}\subset\mathbb{R}^{d}$ from the feasible domain and then receives a stochastic reward $r_{t}\in\mathbb{R}$. We use the notation $\mathcal{X}t$ to indicate that the arm set may vary over time, capturing many practical scenarios where available options change dynamically. For instance, in product recommendation systems, items can be added or removed, requiring the algorithm to adapt accordingly. Besides, we denote the learner’s action by $X_{t}$ to emphasize its stochastic nature, which may depend on past data captured by the filtration $\mathcal{F}_{t}=\sigma(X_{1},r_{1},\dots,X_{t-1},r_{t-1})$. In GLBs, conditioned on $\mathcal{F}_{t}$, the reward $r_{t}$ follows a canonical exponential family distribution with the natural parameter given by the linear model $z_{t}=X_{t}^{\top}\theta_{*}$.
@@
Here, $\theta_{*}\in\mathbb{R}^{d}$ is a $d$-dimensional vector unknown to the learner. The function $h(r,\tau)$ is the base measure, which provides the intrinsic weighting of the variable $r$, while $m(z)$ is twice continuously differentiable function used for normalizing the distribution. Besides, $g:\mathbb{R}\rightarrow\mathbb{R}$ is the dispersion function controlling the variability of the distribution and $\tau\in\mathbb{R}$ is a known parameter. The expectation and variance of the exponential family distribution can be calculated as $\mathbb{E}[r\>|\>z]=\mu(z)\triangleq m^{\prime}(z)$ and $\mbox{Var}(r\>|\>z)=g(\tau)m^{\prime\prime}(z)$ (Wainwright and Jordan, 2008, Proposition 3.1). Common examples of (1) include Gaussian, Bernoulli, and Poisson distributions. The goal of the learner is to maximize the cumulative expected reward, which is equivalent to minimizing the regret,
@@
where $\mathbf{x}_{t,*}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}_{t}}\mu(\mathbf{x}^{\top}\theta_{*})$ is the optimal action. Besides, we have the following standard boundness assumptions used in the GLB literature (Filippi et al., 2010).
@@
The link function $\mu$ is twice differentiable over its feasible domain. Moreover, there exist constants $c_{\mu}>0$ and $L_{\mu}>0$ such that $c_{\mu}\leq\mu^{\prime}(z)\leq L_{\mu}$ for all $z\in[-S,S]$. Consequently, the function $m$ is strictly convex and $\mu$ is strictly increasing.
GLM-UCB Method and Potentially Large Constant. The canonical algorithm for the GLB problem is GLM-UCB (Filippi et al., 2010), which resolves the exploration-exploitation trade-off with an upper‑confidence‑bound strategy (Agrawal, 1995). Under Assumptions 12 and an additional condition that the reward $r_{t}$ is non-negative and almost surely bounded for all $t\in[T]$, GLM-UCB achieves the regret of $\mathcal{O}\big(\kappa(\log T)^{\frac{3}{2}}\sqrt{T}\big)$, where the $\mathcal{O}(\cdot)$-notation is used to highlight the dependence on $\kappa$ and time horizon $T$. The dependence on the horizon $T$ matches that of the linear case, where the $\widetilde{\mathcal{O}}(\sqrt{T})$ rate is nearly optimal (Dani et al., 2008). The bottleneck lies in its linear dependence on the constant $\kappa$, which is defined by
@@ 2.2 New Progress with Self-Concordance
The undesirable linear dependence on $\kappa$ has motivated the development of algorithms with improved theoretical guarantees. By leveraging the self-concordance of the loss, rooted in convex optimization and later used in the analysis of logistic regression (Bach, 2010), recent studies (Russac et al., 2021; Lee et al., 2024; Sawarni et al., 2024) have derived regret bounds with substantially reduced dependence on $\kappa$ for GLB. Following this line, we also adopt the self-concordance assumption here.
@@
Assumption 3 holds for many widely used GLMs. For GLMs where the reward is almost surely bounded in $[0,R]$, the link function satisfies Assumption 3 with coefficient $R$ (Sawarni et al., 2024). For example, the Bernoulli distribution is $1$-self-concordant. Many unbounded GLMs also satisfy self-concordance, including the Gaussian distribution ($R=0$), Poisson distribution ($R=1$), and Exponential distribution ($R=0$). Leveraging self-concordance, Lee et al. (2024) and Sawarni et al. (2024) established improved regret bounds of order $\widetilde{\mathcal{O}}(\sqrt{T/\kappa_{*}})$. In these results, the potentially large constant $\kappa_{*}$ appears at the denominator, which largely improves the $\widetilde{\mathcal{O}}(\kappa\sqrt{T})$ bound by Filippi et al. (2010). However, their methods incur still $\mathcal{O}(t)$ time and space complexities per round. Our goal is to design a method with low computation cost while maintaining strong regret guarantees.
@@
Our GLM assumptions are aligned with the recent work (Lee et al., 2024), which are more general than the canonical GLM formulation introduced in Filippi et al. (2010) and later adopted in Sawarni et al. (2024). Besides Assumptions 1 and 2Filippi et al. (2010) further requires the rewards to be almost surely bounded, which automatically implies self-concordance and thus satisfies Assumption 3 as shown by Sawarni et al. (2024, Lemma 2.2). Beyond bounded distributions, our GLM formulation accommodates unbounded ones, such as Gaussian or Poisson. ◆
@@ 3.1 OFU Principle and Computational Challenge
Computational Challenge. To the best of our knowledge, most existing OFU-based methods for GLBs rely on maximum likelihood estimation (MLE) to estimate $\theta_{*}$ and construct the confidence set. Specifically, the estimator is computed as
@@ 3.2 Jointly Efficient Method
$\widetilde{\ell}_{t}(\theta)\triangleq\langle\nabla\ell_{t}(\theta_{t}),\theta-\theta_{t}\rangle+\frac{1}{2}\lVert\theta-\theta_{t}\rVert^{2}_{\nabla^{2}\ell_{t}(\theta_{t})}~~~\mbox{and}~~~H_{t}\triangleq\lambda I_{d}+\sum_{s=1}^{t-1}\nabla^{2}\ell_{s}(\theta_{s+1}).$ (4)
The above two components play important roles in achieving low computation cost while maintaining statistical efficiency. The loss function $\widetilde{\ell}_{t}(\theta)$ serves as a second-order approximation of the original loss, which preserves the curvature information of the current loss function while ensures that the resulting optimization problem can be solved efficiently. The local matrix can also be expressed as $H_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}\mu^{\prime}(X_{s}^{\top}\theta_{s+1})/g(\tau)\cdot X_{s}X_{s}^{\top}$, where $X_{s}\in\mathbb{R}^{d}$ is the action selected by the learner. The matrix explicitly captures the non-linearity of the link function by $\mu^{\prime}(X_{s}^{\top}\theta_{s+1})$ and retains the curvature information of historical loss functions until time t1t-1. Since the optimization problem (3) is quadratic optimization over an Euclidean ball, it can be solved with a computation cost of $\mathcal{O}(d^{3})$, independent of time $t$. Further details on solving (3) are provided in Appendix A.3.
@@
Let $\delta\in(0,1]$. Set the step size to $\eta=1+RS$ and the regularization parameter to $\lambda=\max\{14d\eta R^{2},6\eta RSL_{\mu}/g(\tau)\}$. For each $t\in[T]$, define the confidence set as
@@
$\beta_{t}(\delta)=\sqrt{\,4\lambda S^{2}+2\eta\ln\!\frac{1}{\delta}+d(6\eta^{2}+\eta)\ln\!\Bigl(1+\frac{L_{\mu}}{\lambda g(\tau)}\Bigr)}=\mathcal{O}\!\left(SR\sqrt{\,d\!\left(S^{2}R+\ln\!\frac{t}{\delta}\right)}\right).$
Then, under Assumptions 1, 2, and 3, we have $\Pr\Bigl[\forall t\geq 1,\;\theta_{*}\in\mathcal{C}_{t}(\delta)\Bigr]\geq 1-\delta$. Besides, the time complexity for solving (3) is $\mathcal{O}(d^{3})$, and the space complexity is $\mathcal{O}(d^{2})$.
Theorem 1 shows that an ellipsoidal confidence set $\mathcal{C}_{t}(\delta)$ can be constructed to quantify the uncertainty of the online estimator $\theta_{t}$ with both statistical and computational efficiency. From a computational perspective, constructing the confidence set relies only on the online estimator, which can be updated with $\mathcal{O}(1)$ time and space complexities. From a statistical view, the radius of the confidence set is independent of $\kappa$, which is crucial for achieving the improved regret bound.
@@
We note that OMD has been used in logistic bandits for constructing confidence sets (Faury et al., 2022; Zhang and Sugiyama, 2023; Lee and Oh, 2024). However, existing methods are not fully jointly efficient compared with our result. Specifically, Faury et al. (2022) required optimization over the original loss at each round, incurring an additional $\mathcal{O}(\log t)$ computation cost and relying on a warm-up strategy to maintain statistical efficiency. In later work, Zhang and Sugiyama (2023) and Lee and Oh (2024) achieved a constant per-round cost but their regret bounds have an extra $\mathcal{O}(\sqrt{\log t})$ multiplicative factor. More importantly, as we will discuss in detail in Section 4, the analyses in these works depend on the specific structure of the logistic model and do not naturally extend to the GLB setting. Our key contribution is to introduce mix-loss-based technique into the confidence set analysis, which not only enables the application of OMD to the broader GLB framework but also improves both statistical and computational efficiency over previous methods for logistic bandits. ◆
Arm Selection and Regret Bound. Based on the ellipsoidal confidence set, one can employ a variety of exploration strategies, not limited to OFU-based methods but also including randomized approaches such as Thompson sampling (Abeille and Lazaric, 2017) for action selection. Here, we adopt the classical OFU-based strategy, where the action $X_{t}$ is selected by solving the bilevel optimization problem $X_{t}=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{X}_{t}}\max_{\theta\in\mathcal{C}_{t}(\delta)}\mu(\mathbf{x}^{\top}\theta).$ Since $\mu$ is an increasing function and $\mathcal{C}_{t}(\delta)$ is an ellipsoid, the OFU-based action selection rule is equivalent to
@@
0: Self-concordant constant $R$, Lipchitz constant $L_{\mu}$, parameter radius $S$, confidence level $\delta$.
@@
Let $\delta\in(0,1]$. Under Assumptions 1, 2, and 3, with probability at least $1-\delta$, Algorithm 1 with parameter $\eta=1+RS$ and $\lambda=\max\{14d\eta R^{2},6\eta RSL_{\mu}/g(\tau)\}$ ensures